The total resistance R of a certain circuit is given by the following formula
R= (1/R1+1/R2)^-1 + R3
R1=10
R2=3
R3=6
Solve for R

Answer by flame8855(424)

R= (1/R1+1/R2)^-1 + R3
R= (1/10 +1/3)^-1+6
R= (13/30)^-1 +6
R= 30/13 + 6 = 108/13